# AnyCap Project: A Unified Framework, Dataset, and Benchmark for Controllable Omni-modal Captioning

### v1

@@ C.3 Qualitative Examples for Main Experiments
C.4 Qualitative Results for Downstream Generation Tasks
@@
To complement the quantitative results reported in the main paper, we provide qualitative examples showcasing the impact of caption refinement by ACM on downstream image and video generation tasks. Specifically, we compare generations produced using initial captions with those using captions refined by ACM. As shown in Fig. 9 and Fig. 10, refined captions lead to outputs that exhibit more accurate visual-semantic correspondence, richer scene details, and improved coherence. These examples further highlight the practical utility of ACM in enhancing multimodal generation quality in real-world applications.
C.5 Human Evaluation Study
Figure 11: Human evaluation of ACM-8B versus GPT-4o with respect to content and style. Bars indicate the proportion of responses rating ACM-8B as worse, similar, or better. Overall, ACM-8B demonstrates superior performance in most cases.
To complement our model-based evaluation, we conducted a human preference study with a dozen evaluators, each holding at least a bachelor’s degree. These evaluators compared the performance of ACM-8B against the base model GPT-4o across both visual and auditory modalities. For each modality, human annotators are presented with outputs from both models and asked to judge which model performs better in terms of Content and Style. The results are summarized in Fig. 11.
Figure 12: Screenshot of human evaluation.

### v2

@@ C.3 Further Results for the Main Experiments
C.4 Downstream Utility
Table 28: Video generation evaluation. Metrics: Visual Quality (VQ), Temporal Consistency (TC), Dynamic Degree ($D$), Text–Video Alignment (TVA), and Factual Consistency (FC).
Model VQ TC $D$ TVA FC Avg
Wan2.1-T2V-14B 2.84 2.78 2.66 2.76 2.77 2.74
 +AnyCap-8B 3.19 (+0.35) 3.06 (+0.28) 3.18 (+0.52) 3.07 (+0.31) 3.00 (+0.23) 3.10 (+0.36)
HunyuanVideo 2.98 2.89 2.94 2.90 2.83 2.91
 +AnyCap-8B 3.61 (+0.63) 3.48 (+0.59) 3.62 (+0.68) 3.34 (+0.44) 3.46 (+0.63) 3.50 (+0.59)
@@ C.4 Downstream Utility
We further evaluate AnyCap’s utility in enhancing noisy captions for downstream video and image generation. Applying AnyCap to refine captions, including those from the Panda video dataset (Wang et al., 2020), we observe that the resulting video generations demonstrate improved visual-semantic grounding and alignment compared to those using the original, unrefined captions (Wan et al., 2025; Kong et al., 2024). Quantitative results for video generation are presented in Table 28. These findings underscore the effectiveness of AnyCap in real-world multimodal generation pipelines.
To complement the quantitative results, we provide qualitative examples showcasing the impact of caption refinement by AnyCap on downstream image and video generation tasks. Specifically, we compare generations produced using initial captions with those using captions refined by AnyCap. As shown in Fig. 10 and Fig. 11, refined captions lead to outputs that exhibit more accurate visual-semantic correspondence, richer scene details, and improved coherence. These examples further highlight the practical utility of AnyCap in enhancing multimodal generation quality in real-world applications.
Appendix D Use of Large Language Models
In preparing this manuscript, we employed advanced language models (e.g., GPT-5, OpenAI, 2025) solely as editorial aids. Their role was limited to refining phrasing, improving readability, and smoothing stylistic inconsistencies across sections. The models were not engaged in developing research questions, proposing methods, analyzing results, or forming conclusions. All substantive ideas, experimental protocols, and technical contributions originated from the authors. Every sentence revised with model assistance was subsequently inspected and approved by human co-authors.